The radii of the top and bottom of a bucket of slant height 13 cm are 9 cm and 4 cm respectively. The height of the bucket is
(a) 10 cm
(b) 12 cm
(c) 15 cm
(d) 16 cm

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Explaination: l² = h² + (R-r)² => 13² = h² + (9 – 4)² ⇒ 169 = h² + 25 ⇒ h² = 169 – 25 ⇒ h² = 144 ⇒ h = 12